# A Survey of Efficient Reasoning for Large Reasoning Models: Language, Multimodality, and Beyond

### v1

@@ A Survey of Efficient Reasoning for Large Reasoning Models:
Xiaoye Qu1 ††thanks: Project Lead and Equal Contributions., Yafu Li1∗, Zhaochen Su2 ††thanks: Core Contributors., Weigao Sun1†, Jianhao Yan3†, Dongrui Liu1†, Ganqu Cui1, Daizong Liu4, Shuxian Liang5, Junxian He6, Peng Li7, Wei Wei8, Jing Shao1, Chaochao Lu1, Yue Zhang3, Xian-Sheng Hua5, Bowen Zhou1,7, Yu Cheng9 ††thanks: Corresponding Author. 1 Shanghai AI Laboratory 2 Soochow University 3 Westlake University 4 Peking University 5 Tongji University 6 The Hong Kong University of Science and Technology 7 Tsinghua University 8 Huazhong University of Science and Technology 9 The Chinese University of Hong Kong
@@ 1 Introduction
Large Language Models (LLMs), such as DeepSeek V3 Liu et al. (2024a), Qwen 2.5 Yang et al. (2024a), LLaMA 3 Dubey et al. (2024), and GPT-4o Hurst et al. (2024), have demonstrated remarkable capabilities across a wide range of tasks Chang et al. (2024); Zhao et al. (2023); Kasneci et al. (2023); Zhu et al. (2024). These models operate in a manner akin to System 1 thinking Frankish (2010); Kahneman (2011); Li et al. (2025e), characterized by fast, intuitive, and automatic decision-making. However, complex reasoning tasks, such as advanced mathematics Lightman et al. (2023); Besta et al. (2024); Ahn et al. (2024) and formal logic Huang and Chang (2022); Pan et al. (2023), demand more deliberate, structured analysis. To tackle these challenges, a new class of models has emerged: Large Reasoning Models (LRMs), including DeepSeek R1 Guo et al. (2025), OpenAI-o1/o3 Jaech et al. (2024), and QwQ Team (2024). These models enhance performance by explicitly generating intermediate reasoning steps, collectively known as chain-of-thought (CoT) Wei et al. (2022), before producing a final answer. Unlike the rapid and heuristic-driven behavior of LLMs, LRMs exhibit deliberate and analytical reasoning, resembling System 2 thinking Evans (2003); Kannengiesser and Gero (2019). This paradigm shift from System 1 to System 2 reasoning Li et al. (2025e) underpins the development of more capable reasoning agents. Yet, this capability comes at a cost: the reasoning process of LRMs tends to be slower and more verbose. As shown in Figure 1, for an elementary-school level math problem, QwQ-32B produces significantly more tokens than its System-1 counterpart, Qwen2.5-32B-Instruct. Moreover, the distributions demonstrate that LRM model QwQ-32B exhibits a significantly greater output length compared to the LLM model Qwen2.5-32B-Instruct. This observation naturally raises a critical question:
@@ 1.2 Position and Contribution
Recently, several survey papers have explored the development of Large Reasoning Models. Besta et al. (2025) and Zeng et al. (2024a) focus on training methodologies for LRMs, while Li et al. (2025e) provide a broader and more comprehensive overview of the field. To offer a more focused perspective, Chen et al. (2025b) examine long chain-of-thought reasoning and categorize existing reasoning paradigms. Additionally, Ji et al. (2025) investigate test-time scaling of LRMs. Despite these valuable contributions, none of the existing surveys specifically address the growing challenge of reasoning efficiency in LRMs, which is an emerging yet crucial topic in the deployment, scalability, and practical application of these models.
For earlier generations of LLMs, Zhou et al. (2024) provide a comprehensive survey of methods for efficient inference. However, efficient reasoning in LRMs introduces a distinct and emerging research challenge, characterized by the generation of excessive and uncontrolled reasoning tokens. While traditional approaches for accelerating inference, e.g., model quantization and pruning Polino et al. (2018); Xiao et al. (2023a); Xia et al. (2023); Wang et al. (2019); Ma et al. (2023); Cheng et al. (2024), and distributed inference systems Patel et al. (2024); Hu et al. (2024a); Zhong et al. (2024); Lin et al. (2024a), can reduce latency and computational costs, our focus in this survey is on improving the efficiency of the reasoning process itself, rather than general inference acceleration. To summarize, this survey makes the following key contributions to the literature:
@@ 2.1 Definition of Reasoning Efficiency
We define reasoning efficiency from a task-distribution perspective, analogous to meta-learning Hospedales et al. (2020). A common goal for LRMs is to reason efficiently across a wide variety of tasks. Let denote an LRM tasked with solving problems. We evaluate its reasoning efficiency over a distribution of tasks , where each task is loosely defined as a pair , with representing a dataset and a corresponding quality metric. We define the reasoning efficiency of model as:
@@ 2.2 Patterns of Reasoning Inefficiency
Despite the effectiveness of LRMs in employing long reasoning chains to address complex problems such as AIME Jaech et al. (2024); DeepSeek-AI (2025); Team (2024), patterns of reasoning inefficiency persist. These inefficiencies manifest as excessive generation of redundant content (e.g., repetitive rephrasing of questions), over-analysis and verification of simple problems (as illustrated in Figure 1), or meandering through superficial and suboptimal thoughts.
@@
Existing LRMs lack explicit optimization for minimizing intermediate reasoning tokens Munkhbat et al. (2025). As a result, these models frequently exhibit redundancy during reasoning Song et al. (2025b), allocating a substantial portion of their output to textual coherence instead of core reasoning advancement Su et al. (2025); Luo et al. (2025a). This type of redundancy increases computational cost in Eq.1, thereby reducing efficiency. For example, in Table 1, the LRM uses 138 tokens simply to rephrase and interpret the question. Verbose explanations and repetitive phrasing further bloat the output without aiding the final solution Munkhbat et al. (2025). Notably, Li et al. (2025b) show that while the overall structure of a chain-of-thought is essential for learning, the content of individual reasoning steps often contributes little to overall model performance and solution accuracy.
@@
Recent studies show that LRMs struggle to allocate their reasoning budget effectively based on task complexity Luo et al. (2025a); Chiang and Lee (2024). For instance, Chen et al. (2025c) observe that models often exhibit unwarranted uncertainty on straightforward queries such as . Rather than producing a concise and direct solution, these models tend to generate multiple redundant reasoning rounds Luo et al. (2025b), exploring unnecessary solution paths. As shown in Figure 1, although the LRM reaches the correct answer in the initial reasoning trace, it performs several additional verification steps, ultimately using nearly forty times as many tokens as a standard instruction-tuned LLM. Such redundancy also leads to increased computational cost in Eq.1 for simple tasks where the model achieves a relatively high quality score , therefore reducing efficiency.
forked edges,
for tree=
child anchor=west,
parent anchor=east,
grow’=east,
anchor=west,
base=left,
font=,
rectangle,
draw=hidden-black,
rounded corners,
align=left,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=3pt,
inner xsep=2pt,
inner ysep=3pt,
line width=0.8pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=12em,font=,,
where level=2text width=12em,font=,,
where level=3text width=12em,font=,,
where level=4text width=10em,font=,,
Efficient Reasoning, ver [
Efficient Reasoning
during Inference (§3), inference
Length Budgeting, inference
TALE (Han et al., 2024),
SketchofThought Aytes et al. (2025),
Planning Tokens (Wang et al., 2023b),
Chain of Draft (Xu et al., 2025a),
S1 (Muennighoff et al., 2025),
SafeChain (Jiang et al., 2025),
DSC (Wang et al., 2024b),
Dynasor (Fu et al., 2024b),
TSP (Wang et al., 2025e),
etc.
, inference, text width=45.6em
System Switch, inference
Dualformer (Su et al., 2024a),
System 1.x Saha et al. (2024),
FaST (Sun et al., 2025a),
HaluSearch (Cheng et al., 2025a),
Dyna-Think (Pan et al., 2024),
etc.
, inference, text width=45.6em
Model Switch, inference
BiLD (Kim et al., 2023),
EAGLE Li et al. (2024e, d),
MEDUSA Cai et al. (2024),
LayerSkip (Elhoushi et al., 2024),
Zooter (Lu et al., 2023),
RouteLLM (Ong et al., 2025),
MixLLM (Wang et al., 2025b),
etc.
, inference, text width=45.6em
Parallel Search, inference
SBoN (Sun et al., 2024a),
TreeBoN Qiu et al. (2024),
STBoN (Wang et al., 2025d),
SelfCalibration (Huang et al., 2025a),
MetaReasoner (Sui et al., 2025),
TPO (Li et al., 2025d),
etc.
, inference, text width=45.6em
Efficient Reasoning
with SFT (§4), sft
Reasoning Chain
Compression, sft
TokenSkip (Xia et al., 2025), SPIRIT-FT Cui et al. (2025), Skip Steps (Liu et al., 2024c),
Distill System 2 (Yu et al., 2024a), C3ot (Kang et al., 2024), Self-Training (Munkhbat et al., 2025), etc.
, sft, text width=45.6em
Latent-Space SFT, sft
Coconut (Hao et al., 2024), CCoT Cheng and Van Durme (2024), CODI (Shen et al., 2025c),
Token Assorted (Su et al., 2025), SoftCoT (Xu et al., 2025b), Heima (Shen et al., 2025a),
Implicit CoT (Deng et al., 2024), LightThinker Zhang et al. (2025a), etc.
, sft, text width=45.6em
Efficient Reasoning
with RL (§5), rl
with Length Reward, rl
O1-Pruner Luo et al. (2025a),
Training Arora and Zanette (2025),
L1 Aggarwal and Welleck (2025),
Kimi-1.5 Team et al. (2025),
DAST (Shen et al., 2025b),
Demisifying Yeo et al. (2025), etc.
, rl, text width=45.6em
without Length Reward, rl
Meta-RL (Qu et al., 2025b),
IBPO (Yu et al., 2025b),
overthink (Chen et al., 2025c), etc.
, rl, text width=45.6em
Efficient Reasoning
during Pre-training (§6), pretrain
Latent-space Pretraining, pretrain
BLT (Pagnoni et al., 2024), LCMs The et al. (2024), CoCoMix (Tack et al., 2025),
LTMs (Kong et al., 2025) , etc.
, pretrain, text width=45.6em
Subquadratic Attention, pretrain
Lightning Attention (Qin et al., 2024e), LASP-2 (Sun et al., 2025b), GLA (Yang et al., 2023),
Gated DeltaNet (Yang et al., 2024d), MoM (Du et al., 2025), Mamba-2 (Dao and Gu, 2024),
RWKV-7 (Peng et al., 2025a), NSA (Yuan et al., 2025a), MoBA (Lu et al., 2025a), etc.
, pretrain, text width=45.6em
Linearization, pretrain
Liger (Lan et al., 2025),
Llamba (Bick et al., 2025),
LoLCATs (Zhang et al., 2024a),
etc.
, pretrain, text width=45.6em
Efficient Reasoning with
Subquadratic Attention, pretrain
TSF (Paliotta et al., 2025), CRQs (Yehudai et al., 2025),
Cosmos-Reason1 (Azzolini et al., 2025),
etc.
, pretrain, text width=45.6em
Future Directions (§7), future
Efficient Multimodality and
Video Reasoning, future
SCoT (Xiang et al., 2025), AL-CoTD (Peng et al., 2025b), Heima (Shen et al., 2025a), etc.
, future, text width=45.6em
Efficient Test-time Scaling
and Infinity Thinking, future
Self-Calibration (Huang et al., 2025a),
Dynamic self-consistency (Wan et al., 2024), etc.
, future, text width=45.6em
Efficient and Trustworthy
Reasoning, future
Deliberative Alignment (Guan et al., 2024), X-Boundary (Lu et al., 2025b), etc.
, future, text width=45.6em
Efficient Reasoning
Applications, future
Chain-of-Retrieval (Wang et al., 2025a),
Overthinking in Agent (Cuadron et al., 2025), etc.
, future, text width=45.6em
Evaluation and Benchmark, future
Overthinking (Chen et al., 2025c), DNA Bench (Hashemi et al., 2025), etc.
, future, text width=45.6em
@@
Wang et al. (2025e) identify a phenomenon termed underthinking, where o1-like LRMs prematurely switch reasoning directions, hindering the development of promising paths. This results in shallow and fragmented reasoning traces, particularly in complex mathematical tasks. Instead of pursuing a coherent, in-depth line of thought, the model hops between multiple approaches superficially, leading to longer reasoning sequences and reduced overall solution quality. Such shallow hopping causes either a reduction in or an elevation of in Eq. 1, thus degrading inference efficiency.
Such reasoning inefficiencies pose significant challenges across training, inference, and real-world applications. Specifically, excessively long CoT sequences hinder reinforcement learning (RL) optimization Yeo et al. (2025); Yu et al. (2025a); Yuan et al. (2025b), leading to instability during RL fine-tuning and excessive memory consumption. Moreover, due to the autoregressive nature of LRM decoding, inference latency increases linearly with reasoning length. This results in high inference costs and degraded user experience, especially when reasoning traces exceed 10,000 tokens. The issue becomes more pronounced in multi-agent systems, where timely generation of plans and responses is critical Huang et al. (2024).
@@ 2.3 Unique Challenges for Efficient Reasoning in the Era of LRMs
One of the fundamental challenges in efficient reasoning is the difficulty of evaluating the utility of each step in a reasoning chain. Unlike classification or regression tasks where loss can be directly computed, it is hard to quantify how much each intermediate reasoning token contributes to the final answer. This lack of granularity makes it difficult to determine which parts of the reasoning can be compressed or pruned Zhang et al. (2025c), and how to balance reasoning conciseness with answer correctness Sprague et al. (2024); Wu et al. (2025b); Yang et al. (2025c). As a result, optimizing for brevity without degrading performance remains a delicate trade-off.
@@
Length control has long been a challenge in previous LLM generations, and it becomes even more critical in the context of reasoning. While token-level constraints can be enforced during decoding, they are often too rigid and oblivious to the semantic structure of reasoning Lee et al. (2025). The ability to “think just enough”, i.e., not too shallowly to miss key logic Wang et al. (2025e), and not too deeply to waste computation Chen et al. (2025c), is essential for reasoning-efficient models, but remains largely unsolved.
@@
Most existing LRMs still rely on the Transformer architecture, which incurs quadratic complexity with respect to input length. This design becomes especially limiting as reasoning traces expand to thousands of tokens or even more. Developing new architectures or efficient approximations that can reason over long contexts without sacrificing performance is a crucial and open direction. Subquadratic attention and linear sequence modeling are promising, but still in their early stages for complex reasoning tasks Paliotta et al. (2025); Yehudai et al. (2025).
@@
Different tasks demand different reasoning depths. Specifically, arithmetic problems may benefit from deep logical traces, while commonsense QA might require only short chains. A single reasoning strategy or length policy often fails to generalize across such varied tasks Sprague et al. (2024). For instance, as shown in Figure 4, the most recent LRM Gemini 2.5 Pro spends thousands of tokens to translate a short Chinese sentence into English. Ensuring efficiency while preserving robustness and adaptability across domains remains an unsolved and nuanced challenge.
@@ 3 Efficient Reasoning during Inference
The computational overhead of LRMs mainly arises from lengthy intermediate reasoning traces. Furthermore, LRMs struggle to reason effectively within a suitable computational budget given the complexity of tasks Chen et al. (2025c); Wu et al. (2025b). To mitigate this challenge, various methods have been proposed to facilitate inference algorithms. These include constraining the reasoning budget, allocating resources across different systems or models, and incorporating parallel search strategies, all aimed at optimizing the trade-off between efficiency and accuracy, as shown in Figure 5. It is important to note that the works discussed in this section are designed to support their proposed inference methods and may involve training models using existing techniques such as supervised fine-tuning (SFT).

### v2

@@ A Survey of Efficient Reasoning for Large Reasoning Models:
Xiaoye Qu1 ††thanks: Project Lead and Equal Contributions., Yafu Li1∗, Zhao-Chen Su2 ††thanks: Core Contributors., Weigao Sun1†, Jianhao Yan3†, Dongrui Liu1†, Ganqu Cui1, Daizong Liu4, Shuxian Liang5, Junxian He6, Peng Li7, Wei Wei8, Jing Shao1, Chaochao Lu1, Yue Zhang3, Xian-Sheng Hua5, Bowen Zhou1,7, Yu Cheng9 ††thanks: Corresponding Author. 1 Shanghai AI Laboratory 2 Soochow University 3 Westlake University 4 Peking University 5 Tongji University 6 The Hong Kong University of Science and Technology 7 Tsinghua University 8 Huazhong University of Science and Technology 9 The Chinese University of Hong Kong
@@ 1 Introduction
Large Language Models (LLMs), such as DeepSeek V3 liu2024deepseek, Qwen 2.5 yang2024qwen2, LLaMA 3 dubey2024llama, and GPT-4o hurst2024gpt, have demonstrated remarkable capabilities across a wide range of tasks chang2024survey; zhao2023survey; kasneci2023chatgpt; zhu2024llama. These models operate in a manner akin to System 1 thinking frankish2010dual; kahneman2011thinking; li202512surveyreasoning, characterized by fast, intuitive, and automatic decision-making. However, complex reasoning tasks, such as advanced mathematics lightman2023let; besta2024graph; ahn2024large and formal logic huang2022towards; pan2023logic, demand more deliberate, structured analysis. To tackle these challenges, a new class of models has emerged: Large Reasoning Models (LRMs), including DeepSeek R1 guo2025deepseek, OpenAI-o1/o3 jaech2024openai, and QwQ qwq-32b-preview. These models enhance performance by explicitly generating intermediate reasoning steps, collectively known as chain-of-thought (CoT) wei2022chain, before producing a final answer. Unlike the rapid and heuristic-driven behavior of LLMs, LRMs exhibit deliberate and analytical reasoning, resembling System 2 thinking evans2003two; kannengiesser2019design. This paradigm shift from System 1 to System 2 reasoning li202512surveyreasoning underpins the development of more capable reasoning agents. Yet, this capability comes at a cost: the reasoning process of LRMs tends to be slower and more verbose. As shown in Figure 1, for an elementary-school level math problem, QwQ-32B produces significantly more tokens than its System-1 counterpart, Qwen2.5-32B-Instruct. Moreover, the distributions demonstrate that LRM model QwQ-32B exhibits a significantly greater output length compared to the LLM model Qwen2.5-32B-Instruct. This observation naturally raises a critical question:
@@ 1.2 Position and Contribution
Recently, several survey papers have explored the development of Large Reasoning Models. besta2025reasoning and zeng2024scaling focus on training methodologies for LRMs, while li202512surveyreasoning provide a broader and more comprehensive overview of the field. To offer a more focused perspective, chen2025reasoningerasurveylong examine long chain-of-thought reasoning and categorize existing reasoning paradigms. Additionally, ji2025test investigate test-time scaling of LRMs. Despite these valuable contributions, none of the existing surveys specifically address the growing challenge of reasoning efficiency in LRMs, which is an emerging yet crucial topic in the deployment, scalability, and practical application of these models.
For earlier generations of LLMs, zhou2024survey provide a comprehensive survey of methods for efficient inference. However, efficient reasoning in LRMs introduces a distinct and emerging research challenge, characterized by the generation of excessive and uncontrolled reasoning tokens. While traditional approaches for accelerating inference, e.g., model quantization and pruning polino2018model; xiao2023smoothquant; xia2023sheared; wang2019structured; ma2023llm; cheng2024survey, and distributed inference systems patel2024splitwise; hu2024inference; zhong2024distserve; lin2024infinite, can reduce latency and computational costs, our focus in this survey is on improving the efficiency of the reasoning process itself, rather than general inference acceleration. To summarize, this survey makes the following key contributions to the literature:
@@ 2.1 Definition of Reasoning Efficiency
We define reasoning efficiency from a task-distribution perspective, analogous to meta-learning hospedales2020metalearningneuralnetworkssurvey. A common goal for LRMs is to reason efficiently across a wide variety of tasks. Let denote an LRM tasked with solving problems. We evaluate its reasoning efficiency over a distribution of tasks , where each task is loosely defined as a pair , with representing a dataset and a corresponding quality metric. We define the reasoning efficiency of model as:
@@ 2.2 Patterns of Reasoning Inefficiency
Despite the effectiveness of LRMs in employing long reasoning chains to address complex problems such as AIME jaech2024openai; deepseekr1; qwq-32b-preview, patterns of reasoning inefficiency persist. These inefficiencies manifest as excessive generation of redundant content (e.g., repetitive rephrasing of questions), over-analysis and verification of simple problems (as illustrated in Figure 1), or meandering through superficial and suboptimal thoughts.
@@
Existing LRMs lack explicit optimization for minimizing intermediate reasoning tokens munkhbat2025selftrainingelicitsconcisereasoning. As a result, these models frequently exhibit redundancy during reasoning song2025prmbench, allocating a substantial portion of their output to textual coherence instead of core reasoning advancement su2025token; luo2025o1. This type of redundancy increases computational cost in Eq.1, thereby reducing efficiency. For example, in Table 1, the LRM uses 138 tokens simply to rephrase and interpret the question. Verbose explanations and repetitive phrasing further bloat the output without aiding the final solution munkhbat2025selftrainingelicitsconcisereasoning. Notably, li2025llms show that while the overall structure of a chain-of-thought is essential for learning, the content of individual reasoning steps often contributes little to overall model performance and solution accuracy.
@@
Recent studies show that LRMs struggle to allocate their reasoning budget effectively based on task complexity luo2025o1; chiang2024over. For instance, chen2025think23overthinkingo1like observe that models often exhibit unwarranted uncertainty on straightforward queries such as . Rather than producing a concise and direct solution, these models tend to generate multiple redundant reasoning rounds luo2025deconstructing, exploring unnecessary solution paths. As shown in Figure 1, although the LRM reaches the correct answer in the initial reasoning trace, it performs several additional verification steps, ultimately using nearly forty times as many tokens as a standard instruction-tuned LLM. Such redundancy also leads to increased computational cost in Eq.1 for simple tasks where the model achieves a relatively high quality score , therefore reducing efficiency.
@@
wang2025thoughts identify a phenomenon termed underthinking, where o1-like LRMs prematurely switch reasoning directions, hindering the development of promising paths. This results in shallow and fragmented reasoning traces, particularly in complex mathematical tasks. Instead of pursuing a coherent, in-depth line of thought, the model hops between multiple approaches superficially, leading to longer reasoning sequences and reduced overall solution quality. Such shallow hopping causes either a reduction in or an elevation of in Eq. 1, thus degrading inference efficiency.
Such reasoning inefficiencies pose significant challenges across training, inference, and real-world applications. Specifically, excessively long CoT sequences hinder reinforcement learning (RL) optimization yeo2025demystifyinglongchainofthoughtreasoning; yu2025dapoopensourcellmreinforcement; yuan2025whatspposcollapselongcot, leading to instability during RL fine-tuning and excessive memory consumption. Moreover, due to the autoregressive nature of LRM decoding, inference latency increases linearly with reasoning length. This results in high inference costs and degraded user experience, especially when reasoning traces exceed 10,000 tokens. The issue becomes more pronounced in multi-agent systems, where timely generation of plans and responses is critical huang2024understanding.
@@ 2.3 Unique Challenges for Efficient Reasoning in the Era of LRMs
One of the fundamental challenges in efficient reasoning is the difficulty of evaluating the utility of each step in a reasoning chain. Unlike classification or regression tasks where loss can be directly computed, it is hard to quantify how much each intermediate reasoning token contributes to the final answer. This lack of granularity makes it difficult to determine which parts of the reasoning can be compressed or pruned zhang2025lessons, and how to balance reasoning conciseness with answer correctness sprague2024cotcotchainofthoughthelps; wu2025lessunderstandingchainofthoughtlength; yang2025towards. As a result, optimizing for brevity without degrading performance remains a delicate trade-off.
@@
Length control has long been a challenge in previous LLM generations, and it becomes even more critical in the context of reasoning. While token-level constraints can be enforced during decoding, they are often too rigid and oblivious to the semantic structure of reasoning lee2025llmscompresschainofthoughttoken. The ability to “think just enough”, i.e., not too shallowly to miss key logic wang2025thoughts, and not too deeply to waste computation chen2025think23overthinkingo1like, is essential for reasoning-efficient models, but remains largely unsolved.
@@
Most existing LRMs still rely on the Transformer architecture, which incurs quadratic complexity with respect to input length. This design becomes especially limiting as reasoning traces expand to thousands of tokens or even more. Developing new architectures or efficient approximations that can reason over long contexts without sacrificing performance is a crucial and open direction. Subquadratic attention and linear sequence modeling are promising, but still in their early stages for complex reasoning tasks paliotta2025thinking; yehudai2025compositionalreasoningtransformersrnns.
@@
Different tasks demand different reasoning depths. Specifically, arithmetic problems may benefit from deep logical traces, while commonsense QA might require only short chains. A single reasoning strategy or length policy often fails to generalize across such varied tasks sprague2024cotcotchainofthoughthelps. For instance, as shown in Figure 4, the most recent LRM Gemini 2.5 Pro spends thousands of tokens to translate a short Chinese sentence into English. Ensuring efficiency while preserving robustness and adaptability across domains remains an unsolved and nuanced challenge.
@@ 3 Efficient Reasoning during Inference
The computational overhead of LRMs mainly arises from lengthy intermediate reasoning traces. Furthermore, LRMs struggle to reason effectively within a suitable computational budget given the complexity of tasks chen2025think23overthinkingo1like; wu2025lessunderstandingchainofthoughtlength. To mitigate this challenge, various methods have been proposed to facilitate inference algorithms. These include constraining the reasoning budget, allocating resources across different systems or models, and incorporating parallel search strategies, all aimed at optimizing the trade-off between efficiency and accuracy, as shown in Figure 5. It is important to note that the works discussed in this section are designed to support their proposed inference methods and may involve training models using existing techniques such as supervised fine-tuning (SFT).